# Online Iterative Reinforcement Learning from Human Feedback with General Preference Model

### v2

@@ B.1 Proof for Theorem 1
$\widetilde{\Gamma}(\pi_{*}^{1},\widetilde{\pi}^{2})$
$=$ $\sup_{P\in\mathcal{P}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\pi_{*}^{1},\widetilde{\pi}^{2})-\hat{P}(x,\pi_{*}^{1},\widetilde{\pi}^{2})]|}{\sqrt{\lambda+\sum_{i=1}^{n}(P(x_{i},a_{i}^{1},a_{i}^{2})-\hat{P}(x_{i},a_{i}^{1},a_{i}^{2}))^{2}}}$
$\leq$ $\sup_{P\in\mathcal{P}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\pi_{*}^{1},\widetilde{\pi}^{2})-\hat{P}(x,\pi_{*}^{1},\widetilde{\pi}^{2})]|}{\sqrt{\lambda-\log(|\mathcal{P}|/\delta)+0.5n\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi_{D}^{1},a^{2}\sim\pi_{D}^{2}}(P(x,a^{1},a^{2})-\hat{P}(x,a^{1},a^{2}))^{2}}}$
$=$ $\sqrt{\frac{2}{n}}\sup_{P\in\mathcal{P}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\pi_{*}^{1},\widetilde{\pi}^{2})-\hat{P}(x,\pi_{*}^{1},\widetilde{\pi}^{2})]|}{\sqrt{\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi_{D}^{1},a^{2}\sim\pi_{D}^{2}}(P(x,a^{1},a^{2})-\hat{P}(x,a^{1},a^{2}))^{2}}}$
@@
B.2 Learning with Pessimism via Uncertainty Bonus
In this subsection, we introduce another offline algorithm, Pessimistic Equilibrium Learning from Human Feedback (PELHF) with Uncertainty Bonus in Algorithm 3. Given an offline dataset $\mathcal{D}_{\mathrm{off}}$, we first obtain the maximum likelihood estimation (MLE) by maximizing Equation (6). Then, we take the lower confidence bound (LCB) for the max-player as the preference estimations by subtracting a bonus function $\beta\Gamma(\cdot,\cdot,\cdot)$:
@@ B.2 Learning with Pessimism via Uncertainty Bonus
2:  Compute the MLE $\hat{P}$ with $\ell_{\mathcal{D}_{\mathrm{off}}}$ defined in Equation (6) and construct bonus as in Equation (18).
@@
where we also set $\beta$ as an upper bound of the $\lambda$-regularized in-sample error. This uncertainty is also characterized by the relative preference function class and shares a similar spirit with the information ratio considered in Zhang, (2023); Ye et al., 2023a ; Ye et al., 2023b , which depicts the uncertainty with respect to the value function class. Also see Definition 3 for a more detailed illustration.
B.2.1 Analysis for Algorithm 3
@@
Comparison between Bonus and Version Space. Compared to the bound in Theorem 1, the bound in Theorem 3 enjoys an additional negative KL divergence term between $\pi_{*}^{1}$ and $\hat{\pi}^{1}$. Both Theorem 1 and Theorem 3 depend on a distribution-shift term between Nash policy $\pi_{*}^{1}$ and the policy $\pi_{D}$ that the data is complied with. The difference is that Theorem 1 enjoys a sharper term $\mathcal{C}$ because of Jensen’s inequality and the expectations are inside the $\sup$ operator. In terms of applicability, the version-space-based pessimism is more preferred because it does not require a point-wise uncertainty estimator, thus applying to general cases. In contrast, point-wise pessimism, or more generally, optimism/pessimism via a biased target may be easier to heuristically approximate in practice, as shown in Coste et al., (2023); Xie et al., (2021); Zhang, (2022); Liu et al., 2023b .
B.3 Analysis for Refined Coverage Condition
In this subsection, we show that with an improved analysis, Algorithm 1 enjoys a refined coverage condition, similar to the coverage notion in Zhang et al., (2023).
@@ B.3 Analysis for Refined Coverage Condition
$J(\pi_{1}^{*},\pi_{2}^{*})-J(\hat{\pi}^{1},\dagger)\leq\min_{\pi^{2}\in\Pi}4\beta\sqrt{\frac{\mathscr{C}((\pi_{*}^{1},\pi^{2}),\pi_{D},\mathcal{P})}{n}}+\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}^{2}}(\pi^{2}),$
@@
$\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}^{2}}(\pi^{2})$ $=\underline{J}(\pi_{*}^{1},\pi^{2})-\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2}).$
@@
We use a different decomposition from the proof of Theorem 1. Recall that
and use the notation
@@
We observe that for any $\pi^{2}$, the following decomposition holds
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\dagger)\leq$ $\underbrace{\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})-J(\hat{\pi}^{1},\dagger)}_{\displaystyle q_{1}}+\underbrace{\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2})-\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})}_{\displaystyle q_{2}}+\underbrace{\underline{J}(\pi_{*}^{1},\pi^{2})-\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2})}_{\displaystyle q_{3}}$
$+$ $\underbrace{J(\pi_{*}^{1},\pi^{2})-\underline{J}(\pi_{*}^{1},\pi^{2})}_{\displaystyle q_{4}}+\underbrace{J(\pi_{*}^{1},\pi_{*}^{2})-J(\pi_{*}^{1},\pi^{2})}_{\displaystyle q_{5}}.$
From the proof of Theorem 1, we know that $q_{1}\leq 0$ and $q_{2}\leq 0$. The term $q_{3}=\underline{J}(\pi_{*}^{1},\pi^{2})-\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2}):=\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}^{2}}(\pi^{2})$ measures the suboptimality gap between $\pi^{2}$ and $\widetilde{\pi}^{2}$ under the pessimistic estimation $\underline{J}$ and Nash policy $\pi_{*}^{1}$. For the term $q_{4}$, we have
@@
Since $\pi_{*}^{2}$ is the best response to $\pi_{*}^{1}$ with respect to $J$, we have $q_{5}\leq 0$. Therefore, we obtain that
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\dagger)\leq 2\beta\widetilde{\Gamma}(\pi_{*}^{1},\pi^{2})+\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}^{2}}(\pi^{2}).$ (19)
Since Equation Equation (19) holds for any $\pi_{2}$, we further have
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\dagger)\leq\min_{\pi^{2}\in\Pi}2\beta\widetilde{\Gamma}(\pi_{*}^{1},\pi^{2})+\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}^{2}}(\pi^{2}).$
@@
  • First, we can simply let $\pi^{2}=\widetilde{\pi}^{2}$, the best response to $\hat{\pi}^{1}$ under the pessimistic estimation, and then the bound becomes $4\beta\sqrt{\mathscr{C}((\pi_{*}^{1},\widetilde{\pi}^{2}),\pi_{D},\mathcal{P})/n}$, which measures the coverage of the dataset $\mathcal{D}$ on $(\pi_{*}^{1},\widetilde{\pi}^{2})$. When the distribution of $\mathcal{D}$ aligns well with the distribution induced by $(\pi_{*}^{1},\widetilde{\pi}^{2})$, the dataset has a good coverage on $(\pi_{*}^{1},\widetilde{\pi}^{2})$ and the term $\mathscr{C}((\pi_{*}^{1},\widetilde{\pi}^{2}),\pi_{D},\mathcal{P})$ becomes small.
  • When $\mathcal{D}$ has a poor coverage on $(\pi_{*}^{1},\widetilde{\pi}^{2})$, i.e., $\mathscr{C}((\pi_{*}^{1},\widetilde{\pi}^{2}),\pi_{D},\mathcal{P})$ is large, our bound adapts to an alternate policy $\pi^{2}$ that achieves a better trade-off between the suboptimality term $\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}^{2}}(\pi^{2})$ and the coverage term $\mathscr{C}((\pi_{*}^{1},\pi^{2}),\pi_{D},\mathcal{P})$. Here the suboptimality term measures the performance gap between $\pi^{2}$ and $\widetilde{\pi}^{2}$ under $\underline{J}(\pi_{*}^{1},\cdot)$. Since $\widetilde{\pi}^{2}$ is the best response to $\hat{\pi}^{1}$ but not necessarily to $\pi_{*}^{1}$, the suboptimality term could be negative, making the summation even smaller.
Comparison to Unilateral Coverage. The unilateral coverage (Cui and Du,, 2022; Zhong et al., 2022a, ) requires the dataset to cover all unilateral pairs $(\pi_{*}^{1},\pi^{2})$ for any $\pi^{2}\in\Pi$, making the bound in Theorem 1 depend on the coverage term of the worst pair. In contrast, the bound in Theorem 4 automatically adapts to the best $\pi^{2}$, achieving the trade-off between the coverage term and the suboptimality term, thus offering a more flexible coverage condition.
Appendix C Proofs for the Online Setting
@@
$J(\pi_{1}^{*},\pi_{2}^{*})-J(\hat{\pi}_{t}^{1},\dagger)$ $=\mathbb{E}_{x\sim d_{0}}\Big{[}J(x,\pi^{*},\pi^{*})-\hat{J}(x,\hat{\pi}_{t}^{1},\widetilde{\pi}_{t}^{2})+\hat{J}(x,\hat{\pi}_{t}^{1},\widetilde{\pi}_{t}^{2})-J(x,\hat{\pi}_{t}^{1},\widetilde{\pi}_{t}^{2})\Big{]}$
@@
where the first inequality uses Equation (20) and $J^{*}(x)=J^{*}(x,\pi^{*},\pi^{*})=0$, in the second inequality we use the definition of $\hat{\pi}_{t}^{2}$, and $\hat{P}(x,\pi,\pi)=0$ in the last equality.
@@
C.1 Uncertainty for the Bradley-Terry Model
@@
This uncertainty bonus is consistent with that of the reward-based RLHF framework up to some multiplicative factor of regularization parameter (Xiong et al.,, 2023) and the boundness parameter.
C.2 Guarantee for Enhancer
@@
3:     Exploitation with the main agent: compute the MLE $\hat{P}_{t}$ with $\ell_{\mathcal{D}_{1:t-1}}$ defined in Equation (6)
@@
where the last inequality uses Equation (20) and Equation (C.2) since $\hat{\pi}_{1}^{t}$ is the Nash equilibrium of $\hat{J}_{t}$. Therefore, we conclude the proof. ∎
@@
Appendix D Technical Lemmas
D.1 Auxiliary Lemmas and Proofs
@@
The existence and uniqueness of the Nash equilibrium is proved in the Proposition 1 in Munos et al., (2023). We proceed to use the uniqueness of the Nash equilibrium and contradiction to prove the lemma. Suppose $\pi_{*}^{1}\neq\pi_{*}^{2}$, since $\pi_{*}^{1}$ is the best response to $\pi_{*}^{2}$ for the max-player, for any $\pi\in\Pi$, we have
@@
Then, we prove that $(\pi_{*}^{2},\pi_{*}^{1})$ is also the Nash equilibrium. Since $P(x,\pi^{1},\pi^{2})=1-P(x,\pi^{2},\pi^{1})$ for any $\pi^{1}$ and $\pi^{2}$, then Equation (24) implies that for any $\pi\in\Pi$,
@@
This demonstrates that $\pi_{*}^{2}$ is the best response to $\pi_{*}^{1}$ for the max-player. Similarly, Equation (23) implies that for any $\pi\in\Pi$,
@@
According to Proposition 1 in Munos et al., (2023), $(\pi_{*}^{1},\pi_{*}^{2})$ is the unique Nash Equilibrium of $\max_{\pi^{1}}\min_{\pi^{2}}J(\pi^{1},\pi^{2})$. According to the definition of saddle point, it sufficies to prove that for any $x\sim d_{0}$,
@@
which contradicts the definition of $\pi_{*}^{1}$. Because of the symmetry of two players, we also get

### v3

@@ C.1 Proof for Theorem 1
$\widetilde{\Gamma}(\pi_{*}^{1},\widetilde{\pi}_{*}^{2})$
$=$ $\sup_{P\in\mathcal{P}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\pi_{*}^{1},\widetilde{\pi}_{*}^{2})-\hat{P}(x,\pi_{*}^{1},\widetilde{\pi}^{2})]|}{\sqrt{\lambda+\sum_{i=1}^{n}(P(x_{i},a_{i}^{1},a_{i}^{2})-\hat{P}(x_{i},a_{i}^{1},a_{i}^{2}))^{2}}}$
$\leq$ $\sup_{P\in\mathcal{P}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\pi_{*}^{1},\widetilde{\pi}_{*}^{2})-\hat{P}(x,\pi_{*}^{1},\widetilde{\pi}_{*}^{2})]|}{\sqrt{\lambda-\log(|\mathcal{P}|/\delta)+0.5n\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi_{D}^{1},a^{2}\sim\pi_{D}^{2}}(P(x,a^{1},a^{2})-\hat{P}(x,a^{1},a^{2}))^{2}}}$
$=$ $\sqrt{\frac{2}{n}}\sup_{P\in\mathcal{P}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\pi_{*}^{1},\widetilde{\pi}_{*}^{2})-\hat{P}(x,\pi_{*}^{1},\widetilde{\pi}_{*}^{2})]|}{\sqrt{\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi_{D}^{1},a^{2}\sim\pi_{D}^{2}}(P(x,a^{1},a^{2})-\hat{P}(x,a^{1},a^{2}))^{2}}}$
@@
C.2 Learning with Pessimism via Uncertainty Bonus
In this subsection, we introduce another offline algorithm, Pessimistic Equilibrium Learning from Human Feedback (PELHF) with Uncertainty Bonus in Algorithm 3. Given an offline dataset $\mathcal{D}_{\mathrm{off}}$, we first obtain the maximum likelihood estimation (MLE) by maximizing (7). Then, we take the lower confidence bound (LCB) for the max-player as the preference estimations by subtracting a bonus function $\beta\Gamma(\cdot,\cdot,\cdot)$:
@@ C.2 Learning with Pessimism via Uncertainty Bonus
2:  Compute the MLE $\hat{P}$ with $\ell_{\mathcal{D}_{\mathrm{off}}}$ defined in (7) and construct bonus as in (19).
@@
where we also set $\beta$ as an upper bound of the $\lambda$-regularized in-sample error. This uncertainty is also characterized by the relative preference function class and shares a similar spirit with the information ratio considered in Zhang [83], Ye et al. [74, 75], which depicts the uncertainty with respect to the value function class. Also see Definition 3 for a more detailed illustration.
C.2.1 Analysis for Algorithm 3
@@
Comparison between Bonus and Version Space. Compared to the bound in Theorem 1, the bound in Theorem 3 enjoys an additional negative KL divergence term between $\pi_{*}^{1}$ and $\hat{\pi}^{1}$. Both Theorem 1 and Theorem 3 depend on a distribution-shift term between Nash policy $\pi_{*}^{1}$ and the policy $\pi_{D}$ that the data is complied with. The difference is that Theorem 1 enjoys a sharper term $\mathcal{C}$ because of Jensen’s inequality and the expectations are inside the $\sup$ operator. In terms of applicability, the version-space-based pessimism is preferred because it does not require a point-wise uncertainty estimator, thus applying to general cases. In contrast, point-wise pessimism, or more generally, optimism/pessimism via a biased target may be easier to heuristically approximate in practice, as shown in Coste et al. [15], Xie et al. [69], Zhang [82], Liu et al. [44].
C.3 Analysis for Refined Coverage Condition
In this subsection, we show that with an improved analysis, Algorithm 1 enjoys a refined coverage condition, similar to the coverage notion in [84].
@@ C.3 Analysis for Refined Coverage Condition
$J(\pi_{1}^{*},\pi_{2}^{*})-J(\hat{\pi}^{1},\dagger)\leq\min_{\pi^{2}\in\Pi}4\beta\sqrt{\frac{\mathscr{C}((\pi_{*}^{1},\pi^{2}),\pi_{D},\mathcal{P})}{n}}+\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}_{*}^{2}}(\pi^{2}),$
@@
$\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}_{*}^{2}}(\pi^{2})$ $=\underline{J}(\pi_{*}^{1},\pi^{2})-\underline{J}(\pi_{*}^{1},\widetilde{\pi}_{*}^{2}).$
@@
Recall that
@@
and $\widetilde{\pi}^{2}_{*}=\min_{\pi^{2}\in\Pi}\underline{J}(\pi^{1}_{*},\pi^{2})$ and $\pi^{\dagger,2}=\min_{\pi^{2}\in\Pi}J(\hat{\pi}^{1},\pi^{2})$. We observe that for any $\pi^{2}$, the following decomposition holds
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\pi^{\dagger,2})\leq$ $\underbrace{J(\pi_{*}^{1},\pi_{*}^{2})-J(\pi_{*}^{1},\pi^{2})}_{\displaystyle q_{1}}+\underbrace{J(\pi_{*}^{1},\pi^{2})-\underline{J}(\pi_{*}^{1},\pi^{2})}_{\displaystyle q_{2}}+\underbrace{\underline{J}(\pi_{*}^{1},\pi^{2})-\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2}_{*})}_{\displaystyle q_{3}}$
$+$ $\underbrace{\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2}_{*})-\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})}_{\displaystyle q_{4}}+\underbrace{\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})-\underline{J}(\hat{\pi}^{1},\pi^{\dagger,2})}_{\displaystyle q_{5}}+\underbrace{\underline{J}(\hat{\pi}^{1},\pi^{\dagger,2})-J(\hat{\pi}^{1},\pi^{\dagger,2})}_{\displaystyle q_{6}}.$
For the term $q_{1}$, since $(\pi_{*}^{1},\pi_{*}^{2})$ is the Nash equilibrium of $J$, we have $q_{1}\leq 0$. By the optimality of $\hat{\pi}^{1}$, term $q_{4}\leq 0$. From the proof of Theorem 1, we know that $q_{5}\leq 0$ and $q_{6}\leq 0$. The term $q_{3}=\underline{J}(\pi_{*}^{1},\pi^{2})-\underline{J}(\pi_{*}^{1},\widetilde{\pi}_{*}^{2}):=\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}_{*}^{2}}(\pi^{2})$ measures the suboptimality gap between $\pi^{2}$ and $\widetilde{\pi}_{*}^{2}$ under the pessimistic estimation $\underline{J}$ and Nash policy $\pi_{*}^{1}$. For the term $q_{2}$, we have
@@
Therefore, we obtain that
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\pi^{\dagger,2})\leq 2\beta\widetilde{\Gamma}(\pi_{*}^{1},\pi^{2})+\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}_{*}^{2}}(\pi^{2}).$ (20)
Since Equation (20) holds for any $\pi_{2}$, we further have
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\pi^{\dagger,2})\leq\min_{\pi^{2}\in\Pi}2\beta\widetilde{\Gamma}(\pi_{*}^{1},\pi^{2})+\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}_{*}^{2}}(\pi^{2}).$
@@
  • First, we can simply let $\pi^{2}=\widetilde{\pi}_{*}^{2}$, the best response to $\pi_{*}^{1}$ under the pessimistic estimation, and then the bound becomes $4\beta\sqrt{\mathscr{C}((\pi_{*}^{1},\widetilde{\pi}_{*}^{2}),\pi_{D},\mathcal{P})/n}$, which measures the coverage of the dataset $\mathcal{D}$ on $(\pi_{*}^{1},\widetilde{\pi}_{*}^{2})$. When the distribution of $\mathcal{D}$ aligns well with the distribution induced by $(\pi_{*}^{1},\widetilde{\pi}_{*}^{2})$, the dataset has a good coverage on $(\pi_{*}^{1},\widetilde{\pi}_{*}^{2})$ and the term $\mathscr{C}((\pi_{*}^{1},\widetilde{\pi}_{*}^{2}),\pi_{D},\mathcal{P})$ becomes small.
  • When $\mathcal{D}$ has a poor coverage on $(\pi_{*}^{1},\widetilde{\pi}_{*}^{2})$, i.e., $\mathscr{C}((\pi_{*}^{1},\widetilde{\pi}_{*}^{2}),\pi_{D},\mathcal{P})$ is large, our bound adapts to an alternate policy $\pi^{2}$ that achieves a better trade-off between the suboptimality term $\mathrm{subopt}^{\pi_{*}^{1},\widetilde{\pi}_{*}^{2}}(\pi^{2})$ and the coverage term $\mathscr{C}((\pi_{*}^{1},\pi^{2}),\pi_{D},\mathcal{P})$. Here the suboptimality term measures the performance gap between $\pi^{2}$ and $\widetilde{\pi}_{*}^{2}$ under $\underline{J}(\pi_{*}^{1},\cdot)$.
Comparison to Unilateral Coverage. The unilateral coverage [17, 86] requires the dataset to cover all unilateral pairs $(\pi_{*}^{1},\pi^{2})$ for any $\pi^{2}\in\Pi$, making the bound in Theorem 1 depend on the coverage term of the worst pair. In contrast, the bound in Theorem 4 automatically adapts to the best $\pi^{2}$, achieving the trade-off between the coverage term and the suboptimality term, thus offering a more flexible coverage condition.
Appendix D Proofs for the Online Setting
@@
$J(\pi_{1}^{*},\pi_{2}^{*})-J(\hat{\pi}_{t}^{1},\dagger)$
$=$ $\mathbb{E}_{x\sim d_{0}}\Big{[}J(x,\pi^{*},\pi^{*})-\hat{J}(x,\hat{\pi}_{t}^{1},\widetilde{\pi}_{t}^{2})+\hat{J}(x,\hat{\pi}_{t}^{1},\widetilde{\pi}_{t}^{2})-J(x,\hat{\pi}_{t}^{1},\widetilde{\pi}_{t}^{2})\Big{]}$
@@
where the first inequality uses (21) and $J^{*}(x)=J^{*}(x,\pi^{*},\pi^{*})=0$, in the second inequality we use the definition of $\hat{\pi}_{t}^{2}$, and $\hat{P}(x,\pi,\pi)=0$ in the last equality.
@@
D.1 Uncertainty for the Bradley-Terry Model
@@
This uncertainty bonus is consistent with that of the reward-based RLHF framework up to some multiplicative factor of regularization parameter [72] and the boundness parameter.
D.2 Guarantee for Enhancer
@@
3:     Exploitation with the main agent: compute the MLE $\hat{P}_{t}$ with $\ell_{\mathcal{D}_{1:t-1}}$ defined in (7)
@@
where the last inequality uses (21) and (D.2) since $\hat{\pi}_{1}^{t}$ is the Nash equilibrium of $\hat{J}_{t}$. Therefore, we conclude the proof. ∎
@@
Appendix E Technical Lemmas
E.1 Auxiliary Lemmas and Proofs
@@
The existence and uniqueness of the Nash equilibrium are proved in Proposition 1 in [46]. We proceed to use the uniqueness of the Nash equilibrium and contradiction to prove the lemma. Suppose $\pi_{*}^{1}\neq\pi_{*}^{2}$, since $\pi_{*}^{1}$ is the best response to $\pi_{*}^{2}$ for the max-player, for any $\pi\in\Pi$, we have
@@
Then, we prove that $(\pi_{*}^{2},\pi_{*}^{1})$ is also the Nash equilibrium. Since $P(x,\pi^{1},\pi^{2})=1-P(x,\pi^{2},\pi^{1})$ for any $\pi^{1}$ and $\pi^{2}$, then (25) implies that for any $\pi\in\Pi$,
@@
This demonstrates that $\pi_{*}^{2}$ is the best response to $\pi_{*}^{1}$ for the max-player. Similarly, (24) implies that for any $\pi\in\Pi$,
@@
According to Proposition 1 in [46], $(\pi_{*}^{1},\pi_{*}^{2})$ is the unique Nash Equilibrium of $\max_{\pi^{1}}\min_{\pi^{2}}J(\pi^{1},\pi^{2})$. According to the definition of the saddle point, it suffices to prove that for any $x\sim d_{0}$,
@@
which contradicts the definition of $\pi_{*}^{1}$. Because of the symmetry of the two players, we also get